# On Experiments

### v1

@@ 7 Deficiency and Quantitative Data Processing Theorems
The converse of the Blackwell-Sherman-Stein theorem states that if $e$ does not divide $e^{\prime}$ then there is a loss function and prior that renders $e^{\prime}$ more useful. The gap in risks is quantified by the deficiency.
@@
We present the proof appearing in [7], with some streamlining.
@@
The proof of the reverse implication of the Blackwell-Sherman-Stein theorem can be recovered by setting $\epsilon=0$. The randomization theorem shows there is a deep connection between differences of risks and deficiency. The following theorem makes this connection precise.
@@
For the proof we require the following simple lemma.
@@
then $\underline{{\mathrm{Risk}}}_{L}^{\pi}(e)\leq\underline{{\mathrm{Risk}}}_{L}^{\pi}(e^{\prime})+\epsilon\left\lVert L\right\rVert_{\infty}$ and $\underline{{\mathrm{Risk}}}_{L}^{\pi}(e^{\prime})\leq\underline{{\mathrm{Risk}}}_{L}^{\pi}(e)+\epsilon\left\lVert L\right\rVert_{\infty}$. By the randomization theorem, this means $\Xi^{\pi}(e,e^{\prime})\leq\epsilon$. This combined with the above lemma completes the proof.
@@
Let $\pi$ be a prior that assigns non zero probability to each unknown. Then $\Xi^{\pi}$ is a metric on experiments modulo equivalence.
Proof  $\Xi^{\pi}$ is obviously non-negative and symmetric. We are required to show that it satisfies the triangle inequality. Let $e,e^{\prime},e^{\prime\prime}$ be experiments, with $T$ and $T^{\prime}$ transitions as in the diagram below.
@@ 7.1 Calculating Deficiency
Experiments $e\in\mathbb{M}(\Theta,\mathcal{Z})$ can be represented by a $|\mathcal{Z}|\times|\Theta|$ column stochastic matrix. Furthermore, the prior distribution $\pi$ can be represented by a vector in $\mathbb{R}^{|\Theta|}$. Using these representations, the directed deficiency can be calculated via linear programming.
@@
Equality is attained in the above if $M_{ij}=\left|\pi_{j}E^{\prime}_{ij}-\pi_{j}\left[FE\right]_{ij}\right|$. Minimizing over $M$ and $F$ yields an optimal solution of $\xi^{\pi}(e,e^{\prime})$.
@@ 8 Conclusion
This paper delivers a powerful, unifying lens on statistical experiments by treating them as morphisms in a category of Markov transitions. It sharpens classic results and opens new avenues:
  • Recasts experiments and decision rules in a single algebraic framework, revealing the essence behind Blackwell’s and Le Cam’s theorems
  • Introduces a general data-processing inequality that extends beyond f-divergences to arbitrary risk functionals
  • Provides a convex-analytic treatment of losses via canonical reparametrization and a neat bias–variance decomposition
  • Turns Le Cam’s abstract deficiency into a concrete linear program, making “how much worse” one experiment is than another computationally tractable

### v2

@@ 7 Deficiency and Quantitative Data Processing Theorems
The converse of the Blackwell-Sherman-Stein theorem states that if $e$ does not divide $e^{\prime}$ then there is a loss function and prior that renders $e^{\prime}$ more useful. The gap in risks is quantified by deficiency.
@@
We present the proof from [17], with some streamlining.
@@
The proof of the reverse implication of the Blackwell-Sherman-Stein theorem can be recovered by setting $\epsilon=0$. The randomization theorem shows that there is a deep connection between differences in risk and deficiency. The following theorem makes this connection precise.
@@
For the proof, we require the following simple lemma.
@@
then $\underline{{\mathrm{Risk}}}_{L}^{\pi}(e)\leq\underline{{\mathrm{Risk}}}_{L}^{\pi}(e^{\prime})+\epsilon\left\lVert L\right\rVert_{\infty}$ and $\underline{{\mathrm{Risk}}}_{L}^{\pi}(e^{\prime})\leq\underline{{\mathrm{Risk}}}_{L}^{\pi}(e)+\epsilon\left\lVert L\right\rVert_{\infty}$. By the randomization theorem, this means $\Xi^{\pi}(e,e^{\prime})\leq\epsilon$. This, combined with the above lemma, completes the proof.
@@
Let $\pi$ be a prior that assigns a nonzero probability to each unknown. Then $\Xi^{\pi}$ is a metric on experiments modulo equivalence.
Proof $\Xi^{\pi}$ is obviously nonnegative and symmetric. We are required to show that it satisfies the triangle inequality. Let $e,e^{\prime},e^{\prime\prime}$ be experiments, with $f$ and $f^{\prime}$ transitions as in the diagram below.
@@ 7.1 Calculating Deficiency
Experiments $e\in\mathbb{M}(\Theta,\mathcal{Z})$ can be represented by a $|\mathcal{Z}|\times|\Theta|$ column stochastic matrix. Furthermore, the prior distribution $\pi$ can be represented by a vector in $\mathbb{R}^{|\Theta|}$. Using these representations, the directed deficiency can be calculated using linear programming.
@@
Equality is attained in the above if $M_{ij}=\left|\pi_{j}E^{\prime}_{ij}-\pi_{j}\left[FE\right]_{ij}\right|$. Minimizing $M$ and $F$ produces an optimal solution of $\xi^{\pi}(e,e^{\prime})$.
@@ 8 Conclusion
This paper delivers a powerful unifying lens on statistical experiments by treating them as morphisms in a category of Markov transitions. It sharpens classic results and opens new avenues:
  • Recasts experiments and decision rules in a single algebraic framework, revealing the essence behind Blackwell’s and Le Cam’s theorems.
  • Introduces a general data processing inequality that extends beyond $\phi$-divergences to arbitrary risk functionals.
  • Provides a convex analytical treatment of losses via canonical reparameterization and neat bias-variance decomposition.
  • Turns Le Cam’s abstract deficiency into a concrete linear program.